# Janus-faces of Temporal Constraint Languages: A dichotomy of expressivity

### v1

@@ 5. Min-clean tuples
Recall the definition of a min-clean tuple: for an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, we say that $t\in R$ is min-clean if $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t_{j})$ for all $i,j\in M(t)$, where
@@
Moreover, a tuple $t\in R$ is min-ready in $R$ if the set $M(t)$ is minimal with respect to inclusion amongst all other sets $M(s)$ for $s\in R$. Clearly, every tuple $t\in R$ that satisfies $|M(t)|=1$ is both min-clean and min-ready in $R$.
@@ 5.1. Min-clean tuples for $\min$
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is smooth and preserved by $\min$, then $E$ contains a min-clean tuple.
@@
If $E$ contains a tuple $t$ with $|M(t)|=1,$ then $t$ is min-clean. Otherwise, take $W$ to be any weakly connected component of $E$. Observe that by connectivity of $E|_{W\times W}$, all vertices $a,b$ contained in $W$ share the same minimal entry. It follows that $\operatorname{\textnormal{minx}}(\min(a,b))=\operatorname{\textnormal{minx}}(a)\cup\operatorname{\textnormal{minx}}(b)$ for all $a,b\in W$. Pick $t^{1},\dots,t^{m}\in E|_{W\times W}$ whose $\sim_{2k}$-classes represent all $\sim_{2k}$-classes appearing in $E|_{W\times W}$. Clearly, this is possible as there are only finitely many $\sim_{2k}$-classes on $\mathbb{Q}^{2k}$. Observe that by smoothness of $E|_{W\times W}$, the set of $\sim_{k}$-classes appearing among $t_{1}^{1},\dots,t_{1}^{m}$ coincides with the set of $\sim_{k}$-classes of $t_{2}^{1},\dots,t_{2}^{m}$. Therefore, the tuple $t\coloneqq{\min^{[m]}}(t^{1},\dots,t^{m})$ is min-clean because it satisfies $\operatorname{\textnormal{minx}}(t_{1})=\bigcup_{a\in W}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(t_{2})$. ∎
As it turns out, in the case of symmetric relations preserved by $\min$, our endeavour to prove the existence of min-clean tuples leads us to the immediate conclusion that every such relation must contain an actual loop. In fact, it even suffices that the relation is cyclic.
@@
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. If $R$ is cyclic and preserved by $\min$, then $R$ contains a loop.
@@
5.2. Min-clean tuples for $\operatorname{mi}$
Lemma 5.3.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{mi}$, $t^{1}\in R$ min-ready, and $t^{2},\dots,t^{m}\in R$. If $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})\neq\varnothing$, $\min(t_{i}^{1})=\dots=\min(t_{i}^{m})$, and $j\in M(t^{1})$, then also $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
Proof.
Let $c=\min(t_{i}^{1})$, and choose mappings $\alpha,\beta$ and $\gamma$ satisfying the requirements in 2.1 in a way such that $\alpha(c)=c$. Consider the tuple $u\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$. Note that $\operatorname{\textnormal{minx}}(u_{i})=\operatorname{\textnormal{minx}}(t_{i}^{1})\cap\dots\cap\operatorname{\textnormal{minx}}(t_{i}^{m})$, and that $\min(u_{i})=c$. By min-readiness of $t^{1}$, also $\min(u_{j})=c$. As $\beta(c),\gamma(c)>\alpha(c)=c$, we must therefore have $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t_{j}^{1})\cap\dots\cap\operatorname{\textnormal{minx}}(t_{j}^{m})\neq\varnothing$. ∎
Lemma 5.4.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. Suppose that $W$ is a weakly connected component of $E$ of pseudo-algebraic length $1$. If $E$ is smooth and preserved by $\operatorname{mi}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)\neq\varnothing$.
Proof.
If $|M(t)|=2$ for all $t\in E$, then by connectivity of $E|_{W\times W}$ we have $\min(a)=\min(b)$ for all vertices $a,b$ appearing in $W$. Take $a,b\in W$ arbitrarily. Since $E$ has pseudo-algebraic length $1$, the $\sim_{k}$-factor of the relation $E|_{W\times W}$ is smooth and $m$-linked for some $m\in\mathbb{N}$. Therefore, there exist $b^{\prime}\sim_{k}b$ and an $m$-fence in $E$ from $a$ to $b^{\prime}$, i.e. an $E$-walk of the form
$(x_{1,0}Ex_{1,1}E\dots Ex_{1,m}=x_{2,m}E^{-1}x_{2,m-1}E^{-1}\dots E^{-1}x_{2,0}=x_{3,0}E\dots E^{-1}x_{n,1}E^{-1}x_{n,0}),$
where $a=x_{1,0}$ and $b^{\prime}=x_{n,0}$. Since $b^{\prime}\sim_{k}b$, we have $\operatorname{\textnormal{minx}}(b)=\operatorname{\textnormal{minx}}(b^{\prime})$. We claim that there is a coordinate on which every lower tip of the above $m$-fence admits its minimal entry, that is
$\operatorname{\textnormal{minx}}(a)\cap\bigcap_{\ell<n}\operatorname{\textnormal{minx}}(x_{\ell,0})\cap\operatorname{\textnormal{minx}}(b^{\prime})\neq\varnothing.$ (7)
Indeed, applying Lemma 5.3 to the tuples $\binom{x_{1,m-1}}{x_{1,m}}$ and $\binom{x_{2,m-1}}{x_{1,m}}$, we obtain $\operatorname{\textnormal{minx}}(x_{1,m-1})\cap\operatorname{\textnormal{minx}}(x_{2,m-1})\neq\varnothing$. Consequently, we also get $\operatorname{\textnormal{minx}}(x_{1,m-2})\cap\operatorname{\textnormal{minx}}(x_{2,m-2})\neq\varnothing$, and, after $m$ steps, $\operatorname{\textnormal{minx}}(a)\cap\operatorname{\textnormal{minx}}(x_{2,0})\neq\varnothing$. We now apply Lemma 5.3 to the edges $\binom{a}{x_{1,1}},\binom{x_{2,0}}{x_{2,1}}$ and $\binom{x_{2,0}}{x_{3,1}}$, and deduce that also $\operatorname{\textnormal{minx}}(x_{1,m})\cap\operatorname{\textnormal{minx}}(x_{3,m})\neq\varnothing$. Continuing like this, we ultimately see that Equation 7 holds true.
Since the $\sim_{k}$-factor of $W$ is finite, we can select finitely many elements $a_{1},a_{2},\dots,a_{\ell}\in W$ such that each $\sim_{k}$-class appearing in $W$ is represented by some $a_{i}$. Moreover, we may choose these elements so that for every $i<\ell$, there exists an $m$-fence in $E$ connecting $a_{i}$ to $a_{i+1}$. By concatenating these fences, we obtain an $m$-fence whose set of lower tips contains all the elements $a_{1},a_{2},\dots,a_{\ell}$. Repeating the argument in the previous paragraph, we see that $\bigcap_{i\leq\ell}\operatorname{\textnormal{minx}}(a_{i})\neq\varnothing$. By the choice of the $a_{i}$, this intersection is equal to $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)$. ∎
Lemma 5.5.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is smooth, has pseudo-algebraic length 1, and is preserved by $\operatorname{mi}$, then $E$ contains a min-clean tuple.
Proof.
Any tuple $t\in E$ with $|M(t)|=1$ is min-clean. Suppose that such a tuple does not exist in $E$, and fix a weakly connected component $W$ of $E$ that has pseudo-algebraic length $1$. It follows that all tuples in $W$ share the same minimal entry $c$. Pick $t^{1},\dots,t^{m}$ such that every $\sim_{2k}$-class of a tuple from $E|_{W\times W}$ is represented by one of $t^{1},\dots,t^{m}$. By smoothness of $E|_{W\times W}$, every $\sim_{k}$-class appearing among $t^{1}_{1},\dots,t^{m}_{1}$ coincides with some $\sim_{k}$-class appearing among $t^{1}_{2},\dots,t^{m}_{2}$, and vice versa. In the definition of $\operatorname{mi}$, choose $\alpha,\beta$ and $\gamma$ such that $\alpha(c)=c$. By Lemma 5.4, we have $\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t^{i}_{j})=\bigcap_{a\in W}\operatorname{\textnormal{minx}}(x)\neq\varnothing$ for $j=1,2.$ The tuple $t\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$ now satisfies $\min(t_{1})=\min(t_{2})=c$, and $\operatorname{\textnormal{minx}}(t_{1})=\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(t_{2})$. ∎
Lemma 5.6.
Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{mi}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
Proof.
Suppose that $t\in R$ satisfies $|M(t)|<n,$ and without loss of generality assume that $1\in M(t)$. By $2$-transitivity of $R$, there are tuples $t^{2},\dots,t^{n}\in R$ such that $1\in M(t^{i})$, $i\notin M(t^{i})$, and $\min(t^{i})=\min(t)$ for all $2\leq i\leq n$. In 2.1, choose $\alpha$ in the definition of $\operatorname{mi}$ such that $\alpha(\min(t))=\min(t).$ It then follows that the tuple $s\coloneqq{\operatorname{mi}^{[n-1]}}(t^{2},\dots,t^{n})$ satisfies $\min(s)=\min(t)$, and $M(s)=\{1\}$. ∎
Lemma 5.7.
Let $R$ be an $n$-ary cyclic and $2$-transitive relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is preserved by $\operatorname{mi}$, then $R$ contains a min-clean tuple.
Proof.
Suppose $\min_{t\in R}|M(t)|>1$, and take $t\in R$ arbitrarily. By Lemma 5.6, $t$ is min-ready and satisfies $\min(t_{1})=\dots=\min(t_{n})=:c$. We claim that $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t_{i})\neq\varnothing.$ Indeed, define $t^{1}\coloneqq t$, and take $t^{2},\dots,t^{n-1}\in R$ to be permutations of $t$ satisfying $t^{i}_{1}=t_{1}$ and $t^{i}_{2}=i+1$ for all $2\leq i\leq n-1$. This is possible by $2$-transitivity of $R$. Moreover, we may choose $t^{n}\in R$ that is a permutation of $t$ and satisfies $t^{n}_{2}=t_{1}$ and $t^{n}_{3}=t_{3}$. Since $\bigcap_{i\leq n-1}\operatorname{\textnormal{minx}}(t_{1}^{i})=\operatorname{\textnormal{minx}}(t_{1})$, Lemma 5.3 applied to the tuples $t^{1},\dots,t^{n-1}$ gives $\bigcap_{i\leq n-1}\operatorname{\textnormal{minx}}(t^{i}_{j})\neq\varnothing$ for all $j\geq 2$. As, by definition, it holds that $\operatorname{\textnormal{minx}}(t^{n}_{3})=\operatorname{\textnormal{minx}}(t^{1}_{3})$ and hence $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t^{i}_{3})\neq\varnothing$, we may apply Lemma 5.3 again to the tuples $t^{1},\dots,t^{n}$. We obtain $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t^{i}_{j})\neq\varnothing$ for all $j\leq n$. By the choice of $t^{i}$, this intersection is for $j=2$ equal to $\bigcap_{i\geq 2}\operatorname{\textnormal{minx}}(t_{i})$.
Let now $s^{1},\dots,s^{n}$ denote all cyclic permutations of the tuple $t$. Since $R$ is cyclic, we have $s^{i}\in R$ for every $i\leq n$. Choose $\alpha$ as in 2.1 such that $\alpha(c)=c,$ and set $s\coloneqq{\operatorname{mi}^{[n]}}(s^{1},\dots,s^{n}).$ Observe that $\min(s)=c$, and $\operatorname{\textnormal{minx}}(s_{i})=\bigcap_{j\leq n}\operatorname{\textnormal{minx}}(t_{j})$ for all $i\leq n$. In particular, $s$ is min-clean. ∎

### v2

@@ 6. Min-clean tuples
Recall the definition of a min-clean tuple: for an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, we say that $t\in R$ is min-clean if $\operatorname{minx}(t_{i})=\operatorname{minx}(t_{j})$ for all $i,j\in M(t)$, where
@@
Moreover, a tuple $t\in R$ is min-ready in $R$ if the set $M(t)$ is minimal with respect to inclusion amongst all other sets $M(s)$ for $s\in R$. Clearly, every tuple $t\in R$ that satisfies $|M(t)|=1$ is both min-clean and min-ready in $R$. Min-ready tuples turn out useful for proving the existence of min-clean tuples.
6.1. Min-clean tuples for $\operatorname{mi}$
The following technical lemma holds true for both digraphs and hypergraphs.
Lemma 6.1.
Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$. If $t^{1}\in R$ is min-ready, and $t^{2},\dots,t^{m}\in R$ satisfy $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{minx}(t^{1}_{i})\cap\dots\cap\operatorname{minx}(t^{m}_{i})\neq\varnothing$, $\min(t_{i}^{1})=\dots=\min(t_{i}^{m})$, and $j\in M(t^{1})$, then also $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{minx}(t^{1}_{j})\cap\dots\cap\operatorname{minx}(t^{m}_{j})\neq\varnothing.$
Proof.
Let $c=\min(t_{i}^{1})$, and choose mappings $\alpha,\beta$ and $\gamma$ satisfying the requirements in Definition 2.1 in a way such that $\alpha(c)=c$. Consider the tuple $u\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$. Note that $\operatorname{minx}(u_{i})=\operatorname{minx}(t_{i}^{1})\cap\dots\cap\operatorname{minx}(t_{i}^{m})$, and that $\min(u_{i})=c$. By min-readiness of $t^{1}$, also $\min(u_{j})=c$. As $\beta(c),\gamma(c)>\alpha(c)=c$, we must therefore have $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{minx}(t_{j}^{1})\cap\dots\cap\operatorname{minx}(t_{j}^{m})\neq\varnothing$. ∎
We now present the core construction for smooth digraphs of pseudo-algebraic length $1$.
Lemma 6.2.
Let $S$ be a binary relation on $\mathbb{Q}^{k}$. If $S$ is smooth, weakly connected, and of pseudo-algebraic length $1$, then either $\langle{S}\rangle_{\operatorname{mi}}$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in\operatorname{supp}{S}}\operatorname{minx}(a)\neq\varnothing$.
Proof.
Let $R\coloneqq\langle{S}\rangle_{\operatorname{mi}}$, and suppose that $|M(t)|=2$ for all $t\in R$. In particular, all elements of $R$ are min-ready. Moreover, observe that $R$ is weakly connected. It follows that $\min(t_{1})=\min(t_{2})$ for all $t\in R$. We may therefore apply Lemma 6.1 to any tuples $t^{1},\dots,t^{m}\in R$ satisfying either $\operatorname{minx}(t^{1}_{1})\cap\dots\cap\operatorname{minx}(t^{m}_{1})\neq\varnothing$ or $\operatorname{minx}(t^{1}_{2})\cap\dots\cap\operatorname{minx}(t^{m}_{2})\neq\varnothing$.
Now, take $a,b\in\operatorname{supp}(S)\subseteq\operatorname{supp}(R)$ arbitrarily. Since $R$ is smooth, weakly connected, and has pseudo-algebraic length $1$, its $\sim_{k}$-factor is $m$-linked for some $m\in\mathbb{N}$. Therefore, there exist $b^{\prime}\sim_{k}b$ and an $m$-fence in $R$ from $a$ to $b^{\prime}$, i.e. an $R$-walk of the form
$(x_{1,0}Rx_{1,1}R\dots Rx_{1,m}=x_{2,m}R^{-1}x_{2,m-1}R^{-1}\dots R^{-1}x_{2,0}=x_{3,0}R\dots R^{-1}x_{n,1}R^{-1}x_{n,0}),$
where $a=x_{1,0}$ and $b^{\prime}=x_{n,0}$. Since $b^{\prime}\sim_{k}b$, we have $\operatorname{minx}(b)=\operatorname{minx}(b^{\prime})$. We claim that there is a coordinate on which every lower tip of the above $m$-fence admits its minimal entry, that is
$\operatorname{minx}(a)\cap\bigcap_{1<\ell<n}\operatorname{minx}(x_{\ell,0})\cap\operatorname{minx}(b^{\prime})\neq\varnothing.$ (7)
Indeed, applying Lemma 6.1 to the tuples $\binom{x_{1,m-1}}{x_{1,m}}$ and $\binom{x_{2,m-1}}{x_{1,m}}$, we obtain $\operatorname{minx}(x_{1,m-1})\cap\operatorname{minx}(x_{2,m-1})\neq\varnothing$. Consequently, we also get $\operatorname{minx}(x_{1,m-2})\cap\operatorname{minx}(x_{2,m-2})\neq\varnothing$, and, after $m$ steps, $\operatorname{minx}(a)\cap\operatorname{minx}(x_{2,0})\neq\varnothing$. We now apply Lemma 6.1 to the tuples $\binom{a}{x_{1,1}},\binom{x_{2,0}}{x_{2,1}}$ and $\binom{x_{2,0}}{x_{3,1}}$ of $R$, and deduce that also $\operatorname{minx}(x_{1,m})\cap\operatorname{minx}(x_{3,m})\neq\varnothing$. Continuing like this, we ultimately see that Equation 7 holds true.
Since the $\sim_{k}$-factor of $S$ is finite, we can select finitely many elements $a_{1},a_{2},\dots,a_{\ell}\in\operatorname{supp}(E)$ such that each $\sim_{k}$-class appearing in $\operatorname{supp}(S)$ is represented by some $a_{i}$. Moreover, we may choose these elements so that for every $i<\ell$, there exists an $m$-fence in $R$ connecting $a_{i}$ to $a_{i+1}$. By concatenating these fences, we obtain an $m$-fence whose set of lower tips contains all the elements $a_{1},a_{2},\dots,a_{\ell}$. Repeating the argument in the previous paragraph, we see that $\bigcap_{i\leq\ell}\operatorname{minx}(a_{i})\neq\varnothing$. By the choice of the $a_{i}$, this intersection is equal to $\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)$. ∎
The existence of min-clean tuples is an almost immediate consequence of Lemma 6.2.
Lemma 6.3.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$. If $E$ is smooth and has pseudo-algebraic length 1, then $E$ contains a min-clean tuple.
Proof.
Any tuple $t\in E$ with $|M(t)|=1$ is min-clean. Suppose that such a tuple does not exist in $E$. Since $E$ is smooth and has pseudo-algebraic length $1$, we can pick $S\subseteq E$ that is smooth, weakly connected, and has pseudo-algebraic length $1$. It follows that all tuples in $\operatorname{supp}(S)$ share the same minimal entry $c$. Pick $t^{1},\dots,t^{m}$ such that every $\sim_{2k}$-class of a tuple from $S$ is represented by one of $t^{1},\dots,t^{m}$. By smoothness of $S$, every $\sim_{k}$-class appearing among $t^{1}_{1},\dots,t^{m}_{1}$ coincides with some $\sim_{k}$-class appearing among $t^{1}_{2},\dots,t^{m}_{2}$, and vice versa. In the definition of $\operatorname{mi}$, choose $\alpha,\beta$ and $\gamma$ such that $\alpha(c)=c$. Applying Lemma 6.2 to $S$, we see that $\bigcap_{i\in[m]}\operatorname{minx}(t^{i}_{j})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(x)\neq\varnothing$ for $j=1,2.$ The tuple $t\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$ now satisfies $\min(t_{1})=\min(t_{2})=c$, and $\operatorname{minx}(t_{1})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)=\operatorname{minx}(t_{2})$. ∎
The following lemma, in the setting of $2$-transitive higher-arity relations $R$, either directly yields a min-clean tuple, or implies that $R$ consists entirely of min-ready tuples.
Lemma 6.4.
Let $n\geq 3$, and let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
Proof.
Suppose that $t\in R$ satisfies $|M(t)|<n,$ and without loss of generality assume that $1\in M(t)$. By $2$-transitivity of $R$, there are tuples $t^{2},\dots,t^{n}\in R$ such that $1\in M(t^{i})$, $i\notin M(t^{i})$, and $\min(t^{i})=\min(t)$ for all $2\leq i\leq n$. In Definition 2.1, choose $\alpha$ in the definition of $\operatorname{mi}$ such that $\alpha(\min(t))=\min(t).$ It then follows that the tuple $s\coloneqq{\operatorname{mi}^{[n-1]}}(t^{2},\dots,t^{n})$ satisfies $\min(s)=\min(t)$, and $M(s)=\{1\}$. ∎
Finally, we derive min-clean tuples for $2$-transitive relations. We point out that the reasoning in the first paragraph of the proof crucially relies on the assumption $n\geq 3$.
Lemma 6.5.
Let $n\geq 3$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$. If $R$ is cyclic and $2$-transitive, then $R$ contains a min-clean tuple.
Proof.
Suppose $\min_{t\in R}|M(t)|>1$, and take $t\in R$ arbitrarily. By Lemma 6.4, $t$ is in particular min-ready, and it satisfies $\min(t_{1})=\dots=\min(t_{n})=:c$. We claim that $\bigcap_{i\leq n}\operatorname{minx}(t_{i})\neq\varnothing.$ Indeed, define $t^{1}\coloneqq t$, and take $t^{2},\dots,t^{n-1}\in R$ to be permutations of $t$ satisfying $t^{i}_{1}=t_{1}$ and $t^{i}_{2}=i+1$ for all $2\leq i\leq n-1$. This is possible by $2$-transitivity of $R$. Moreover, we may choose $t^{n}\in R$ that is a permutation of $t$ and satisfies $t^{n}_{2}=t_{1}$ and $t^{n}_{3}=t_{3}$. Since $\bigcap_{i\leq n-1}\operatorname{minx}(t_{1}^{i})=\operatorname{minx}(t_{1})$, Lemma 6.1 applied to the tuples $t^{1},\dots,t^{n-1}$ gives $\bigcap_{i\leq n-1}\operatorname{minx}(t^{i}_{j})\neq\varnothing$ for all $j\geq 2$. As, by definition, it holds that $\operatorname{minx}(t^{n}_{3})=\operatorname{minx}(t^{1}_{3})$ and hence $\bigcap_{i\leq n}\operatorname{minx}(t^{i}_{3})\neq\varnothing$, we may apply Lemma 6.1 again to the tuples $t^{1},\dots,t^{n}$. We obtain $\bigcap_{i\leq n}\operatorname{minx}(t^{i}_{j})\neq\varnothing$ for all $j\leq n$. By the choice of $t^{i}$, this intersection is for $j=2$ equal to $\bigcap_{i\geq 2}\operatorname{minx}(t_{i})$.
Let now $s^{1},\dots,s^{n}$ denote all cyclic permutations of the tuple $t$. Since $R$ is cyclic, we have $s^{i}\in R$ for every $i\leq n$. Choose $\alpha$ as in Definition 2.1 such that $\alpha(c)=c,$ and set $s\coloneqq{\operatorname{mi}^{[n]}}(s^{1},\dots,s^{n}).$ Observe that $\min(s)=c$, and $\operatorname{minx}(s_{i})=\bigcap_{j\leq n}\operatorname{minx}(t_{j})$ for all $i\leq n$. In particular, $s$ is min-clean. ∎
@@ 6.2. Min-clean tuples for $\min$
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\min$. If $E$ is smooth, then $E$ contains a min-clean tuple.
@@
If $E$ contains a tuple $t$ with $|M(t)|=1,$ then $t$ is min-clean. Otherwise, take $S\subseteq E$ to be smooth and weakly connected. Observe that by connectivity of $S$, all vertices $a,b$ contained in $\operatorname{supp}(S)$ share the same minimal entry. It follows that $\operatorname{minx}(\min(a,b))=\operatorname{minx}(a)\cup\operatorname{minx}(b)$ for all $a,b\in\operatorname{supp}(S)$. Pick $t^{1},\dots,t^{m}\in S$ whose $\sim_{2k}$-classes represent all $\sim_{2k}$-classes appearing in $S$. Clearly, this is possible as there are only finitely many $\sim_{2k}$-classes on $\mathbb{Q}^{2k}$. Observe that by smoothness of $S$, the set of $\sim_{k}$-classes appearing among $t_{1}^{1},\dots,t_{1}^{m}$ coincides with the set of $\sim_{k}$-classes of $t_{2}^{1},\dots,t_{2}^{m}$. Therefore, the tuple $t\coloneqq{\min^{[m]}}(t^{1},\dots,t^{m})$ is min-clean because it satisfies $\operatorname{minx}(t_{1})=\bigcup_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)=\operatorname{minx}(t_{2})$. ∎
As it turns out, in the case of cyclic relations preserved by $\min$, our endeavour to prove the existence of min-clean tuples leads us to the immediate conclusion that every such relation must contain an actual loop.
@@
Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\min$. If $R$ is cyclic, then $R$ contains a loop.